# Frontier AI’s Impact on the Cybersecurity Landscape

### v3

@@ 4.5 Defense For Humans
LLMs show promise in detecting social bots [74], and AI has been deployed for real-world fraud detection [131, 268, 60]. Research demonstrates multi-modal agents capable of detecting phishing websites [31]. AI-driven solutions are increasingly being applied to real-world scenarios. For example, one company deployed an AI-based honeypot to combat vishing, using a chatbot with a “grandma” voice to occupy fraudsters and waste their time [158]. AI-based deepfake detection continues as an active research area [187], with exploration of LLMs’ capabilities in identifying fake images [5]. Another significant direction focuses on AI-powered misinformation detection [12, 61, 142], often employing LLMs to identify LLM-generated content. Watermarking techniques represent another promising detection approach [277, 174, 141, 278, 87].
@@
Despite these active research efforts, Figure 2 shows current defensive techniques struggling to match sophisticated attacks. Research reveals significant limitations in AI systems’ defensive capabilities. For example, studies examining GPT-4’s ability to detect fake images [5] conclude the model lacks reliability and shows limitations with complex tasks. While many papers propose AI-powered defenses against deepfakes and social bots, most address traditional attack scenarios rather than novel or sophisticated attacks enabled by frontier AI. We need studies on adversarial dynamics between defense-based and attack-based AI systems—for example, how defensive AI can detect deepfakes or social bots generated by an AI attacker. This gap underscores the urgent need for research to develop AI-driven defenses against AI-driven human-targeted attacks.
@@ 5 Expert Survey
Expert surveys are a widely used approach for forecasting AI development timelines and risks [84, 217]. Accordingly, we conduct an expert survey to assess experts’ perceptions of frontier AI’s current impact on cybersecurity and their forecasts of its future impact.
@@ 6.2 Asymmetry-Based Prediction
Asymmetry in cost of failures. A clear asymmetry exists between attackers and defenders in their goals and requirements. Attackers need just one successful exploit, while defenders must guard against all possible threats and vulnerabilities [212]. This fundamental difference severely constrains defenders’ failure tolerance. For AI-driven threat detection systems, both false positives and false negatives carry serious consequences: false positives disrupt normal functionalities, while false negatives leave threats uncovered. Even a single false negative can compromise the entire defense. Similarly, for automatic patching, ensuring accuracy before deploying machine-generated code is essential, raising capability requirements and necessitating additional verification processes that inevitably require human intervention. Compounding this issue, frontier AI’s inherent probabilistic nature introduces errors that further restrict its deployment in critical defense systems.
@@
Different priorities of scalability and reliability. Attackers prioritize scalability, while defenders emphasize reliability. This makes AI a natural choice for attackers, as its primary capability is reducing human involvement and increasing automation. AI lowers barriers for attackers by decreasing requirements for deep expertise, specific target system knowledge, and execution time and effort. Even if an attack fails against one system, it may succeed against another. For example, in large-scale attacks, attackers may simultaneously target thousands or millions of systems (e.g., Mirai Botnet [44]). Conversely, frontier AI models face reliability limitations in robustness, transparency, and generalizability [279, 222, 86]. When used for malware detection, users cannot understand why specific software is flagged as malicious or predict when false predictions will occur. This reliability deficit hinders defenders from fully trusting and deploying AI techniques in critical defense systems including decision-making processes. This asymmetry may be magnified in critical physical infrastructure, where attackers exploit these infrastructures and propagate the damage to other systems.
@@ 6.3 Evolutionary Attack-Defense Dynamics
Long-term, however, the evolutionary trajectory may gradually shift toward defenders as several counterfactors emerge. First, defenders could gain priority access to advanced techniques (e.g., OpenAI’s early access for safety researchers [170]). As frontier AI advances, defenders benefit from improved automation, transitioning from manual patch development to AI-generated patches with automated verification. Second, as defenders use AI to systematically discover and patch vulnerabilities, systems become increasingly robust. In particular, by leveraging AI for formal verification and adopting a secure-by-design approach, defenders can build fundamentally more secure systems with provable guarantees. This continuous security improvement creates an environment where finding new vulnerabilities becomes significantly more challenging and costly for attackers. Once defenders achieve stronger system security, attackers must invest substantially more resources to find attack vectors, potentially making many attacks impractical or cost-prohibitive. Lastly, as regulatory frameworks and law enforcement mechanisms mature, attackers can face more severe penalties once caught, substantially raising failure costs and potentially deterring malicious activities.
@@ 7.1 Continuous Marginal Risk Assessment
Build benchmarks and red-teaming approaches for new risks in AI agents. We need to conduct fine-grained risk classification for agents and build specialized benchmarks for each category, given that agents are highly specialized (e.g., web agents, coding agents, and personal assistants). We can also classify risks based on their targets: AI components or symbolic components. Currently, the predominant risk is through indirect prompt injection, attacking AI through symbolic components (tools). However, there are only a few early attacks targeting symbolic components through AI, such as AI-generated exploits against coding agents. This underscores the critical need for thorough risk evaluations of agents. Existing red-teaming for agents only explores limited risks of specific agents (e.g., interpreter abuse of code agents [236], prompt injection of web agents [137, 69], indirect prompt injection [243], and taint-style vulnerability [143] of various agents). Developing automated red-teaming frameworks for broader agentic risks is necessary, such as assessing multi-agent system risks through game theory or designing new attacks for systems with multi-modal agents.
Continuously ensure quality and realism of benchmarks. If a security benchmark is flawed, agents can “cheat” by completing the task via unintended pathways rather than the intended exploit. Given the complexity of these tasks—often spanning multiple files, services, and hosts—benchmarks must be carefully constructed, validated, and demonstrably solvable. For instance, Transluce noted that several of the Intercode-CTF benchmark tasks are broken or unsolvable [154, 257]. Moreover, human expertise remains essential for creating quality environments and data. Research shows that the automated collection of vulnerable code commits without human oversight often yields non-vulnerable or security-unrelated samples [228]. Balancing quality with scale requires combining expert-created seed samples with automated methods [260]. To ensure robust assessment quality, a combination of LLM judges, dynamic evaluation methods, and static program analysis techniques is recommended. Benchmarks should reflect realistic scenarios; real-world testing would be ideal, but raises challenges like verification difficulties and legal/ethical constraints. A practical approach is to replicate production systems in controlled environments with built-in verification mechanisms to assess frontier AI capabilities. Finally, cybersecurity benchmarks must be continuously updated to keep pace with the rapidly evolving threat landscape. Timely updates surface emerging AI-enabled capabilities, allowing practitioners to prioritize mitigations and validate defenses before widespread exploitation.
@@ 7.2 Enhance Cyber Defenses
Enhance attack detection. We first recommend constructing high-quality benchmarks that minimize the critical data and label issues, such as duplicated data, spurious correlations, biased samples, and wrong and missing labels. A recent research [15] provided a comprehensive summary of these issues. Given the challenges of collecting high-quality real-world attack data, collecting data from controlled environments [24] and creating synthetic data [247] could be efficient ways. Constructing hidden test sets with unseen attacks will help continuously retrain AI-driven defenses; once the currently hidden test sets are used for training, new test sets must be constructed and kept separate from the training data. Such practices have been applied in the natural language field: hiding test sets (e.g., SQuAD [186]), updating test sets with recent data (e.g., LiveBench [244]), or estimating and resolving training-test overlap [271].
@@
Improve the automation of defenses with provable guarantees. We believe frontier AI will significantly advance the scalability and automation of formal verification and call for increased efforts in this direction. First, we can use and fine-tune models to generate more complex proofs and constraints for theorem proving, as well as invariants for complex programs [258]. Second, we can use LLMs to generate necessary inputs for formal verification, such as program invariants and function summaries, where enabling verifiable generation with correctness guarantees is crucial. More broadly, secure code generation is a promising direction that can be used to turn unsafe latency code into safe counterparts. Early research has explored strategies like prefix learning [100] and constrained decoding [76]. We suggest focusing more on fundamental improvements in the models or agents, such as RL-based fine-tuning to improve their inherent capability of generating safe code. We can also design customized models and agents to transpile legacy programs into newer and memory-safe languages (e.g., C to Rust [67]). Third, AI can also improve constrained solvers. For example, SOTA solvers rely on random/rule-based mechanisms for backtracking and re-selecting verification paths to resolve conflicts. With frontier AI, solvers could learn more efficient strategies for making these critical decisions. Alongside building better AI-enhanced defenses, society must improve overall security posture and actively deploy these new defensive technologies.
@@
Lesson and recommendations based on AIxCC. The AIxCC competition [1] highlights valuable insights into the current capabilities and limitations of AI-powered cybersecurity systems. The competition required teams to build systems to automatically analyze large codebases and generate PoCs and patches. Most systems in the final still heavily rely on traditional techniques for vulnerability detection and PoC generation. It shows that AI capabilities in these fundamental steps require significant improvement, which aligns with our study. In contrast, patch generation has shown significant evolution with heavy reliance on LLMs, which represent the SOTA in automated code generation. However, these systems still incorporate substantial human knowledge through carefully crafted workflows and prompts. Notably, team Theori [249] adopts a more agentic strategy by using a large number of agents across all steps, enabling them to achieve third place with simpler code structure and reduced human efforts.  These observations align with our recommendations on developing more advanced agents with agent-determined workflows and more program analysis tools, where agent-determined workflows can improve the generalizability and domain-specific tools can improve the agent’s capabilities in security tasks (e.g., static analysis-based context retrieval and fuzzing-based patch validation). In addition, the agentic systems can provide environment, data, and reward signals to continuously improve models’ capabilities in such systems.

### v4

@@ 4.5 Defense For Humans
LLMs show promise in detecting social bots [ferrara2023social], and AI has been deployed for real-world fraud detection [nvdia, zanke2023ai, dhieb2020secure]. Research demonstrates multi-modal agents capable of detecting phishing websites [cao2024phishagent]. AI-driven solutions are increasingly being applied to real-world scenarios. For example, one company deployed an AI-based honeypot to combat vishing, using a chatbot with a “grandma” voice to occupy fraudsters and waste their time [granny]. AI-based deepfake detection continues as an active research area [rana2022deepfake], with exploration of LLMs’ capabilities in identifying fake images [al2023gpt]. Another significant direction focuses on AI-powered misinformation detection [anirudh2023multilingual, dhiman2024gbert, liu2024preventing], often employing LLMs to identify LLM-generated content. Watermarking techniques represent another promising detection approach [zhang2024remark, pan2024markllm, liu2024survey, zhao2024sok, gunn2024undetectable].
@@
Despite these active research efforts, Figure 2 shows current defensive techniques struggling to match sophisticated attacks. Research reveals significant limitations in AI systems’ defensive capabilities. For example, studies examining GPT-4’s ability to detect fake images [al2023gpt] conclude the model lacks reliability and shows limitations with complex tasks. While many papers propose AI-powered defenses against deepfakes and social bots, most address traditional attack scenarios rather than novel or sophisticated attacks enabled by frontier AI. We need studies on adversarial dynamics between defense-based and attack-based AI systems—for example, how defensive AI can detect deepfakes or social bots generated by an AI attacker. This gap underscores the urgent need for research to develop AI-driven defenses against AI-driven human-targeted attacks.
5 Empirical Evaluation
In Tables 6 and 7, we summarize the best-performing models and agents on representative attack and defense benchmarks discussed in Table 1. The results align with our earlier observations: current AI agents are more frequently applied to offensive tasks than defensive ones. Moreover, existing agents tend to use a limited set of domain-specific tools, with limited integration of advanced program analysis techniques that are useful for defensive tasks. In this section, we conduct an empirical evaluation of widely used agents on selected benchmarks to further quantify the performance of off-the-shelf agents in security analysis.
Here, we first summarize the empirical evaluation results. On the attack side, the agent demonstrates stronger capabilities in the earlier stages than in the final exploit step. It also occasionally breaks the environments it constructed. On the defense side, the agent struggles to generate PoCs for large codebases due to limited code analysis capabilities and the lack of domain-specific tools (Figure 5). Once a PoC is generated, security patches require three steps, root cause analysis, patch generation, and patch validations. Root causes of most security vulnerabilities are straightforward with sanitizer reports and stack traces. However, AI agents struggle with the cases where the root cause is not located in the stack traces. Since most security patches are self-contained [li2017large, mei2024arvo], patch generation typically involves light-weighted modifications, and generating additional test cases for patch validation is generally less critical. Accordingly, we observe that even without provided unit tests, AI agents can still achieve a high patch success rate by directly applying their generated patches without creating additional functional tests. This is different from fixing general software engineering bugs (e.g., SWE-bench [jimenez2023swe]), where AI agents need to generate thorough test cases because patches typically modify the program’s logic.
5.1 Offensive Capability Evaluation
Experiment Setup and Design. We choose AutoPenBench [gioacchini2024autopenbench] for this experiment. As the SOTA penetration testing benchmark, AutoPenBench includes end-to-end attacks covering the most comprehensive attack steps and designs its attack tasks based on real-world systems. We use the OpenHands [wang2024openhands] agent with Claude Sonnet 4.5 as the core LLM, as OpenHands is a general agent scaffold, and this combination reports the SOTA performance on existing cybersecurity benchmarks (e.g., CyberGym [wang2025cybergym] and Cybench [zhang2024cybench]. Note that we further equip the OpenHands agent with the kali environment provided by the AutoPenBench, enabling it to use domain-specific tools supported by kali. We denote the agent as “OpenHands-Pen”.
AutoPenBench includes both in-vitro and real-world evaluations. The in-vitro suite comprises 22 synthetic tasks across four categories—Access Control (AC), Web Security (WS), Network Security (NS), and Cryptography (CRPT)—while the real-world suite covers 11 publicly disclosed CVEs. In this experiment, we run OpenHands-Pen over all tasks and report the attack success rate and average cost for each task.
Table 3: Offensive evaluation results of OpenHands-Pen on AutoPenBench. The table reports the Success Rate (SR), average number of LLM queries, and average cost per task.
  Tasks SR Avg Queries Avg Cost (\$)
In-vitro 22 0.59 20.1 0.139
Real-world 11 0.55 50.0 0.454
Total 33 0.58 - 0.202
Figure 3: Success Rate of each attack step for real-world tasks.
Figure 4: Top 10 commands executed by OpenHands-Pen.
Figure 5: Top 10 commands executed by OpenHands-Def.
Experiment Results. Table 3 summarizes the evaluation results of OpenHands-Pen on both in-vitro and real-world tasks. The agent achieves an overall success rate of 0.58, nearly matching the human-assisted agent with GPT-4o reported in AutoPenBench (0.64), demonstrating the advancement of models. Besides, we observe that although the success rate is similar, OpenHands-Pen used 2-3x more queries to solve the real-world tasks, indicating that they are more challenging than synthetic ones. Figure 3 shows the per-step attack success rate, where the overall success rate is the same as the exploitation rate. Here, later attack steps (delivery and exploitation) have a lower attack success rate compared to earlier ones, which is aligned with their corresponding difficulty. Figure 4 further reports the top-10 frequent tools used by the agent, which shows that aside from Metasploit [metasploit_framework], OpenHands-Pen still mostly uses the common bash commands. The observation is aligned with the defensive evaluation in Section 5.2.
More specifically, the agent exhibits different performance across different types of in-vitro tasks. On access control tasks, it achieves a 0.60 success rate by attempting common weak credentials rather than using brute-force tools like Hydra. On web security tasks, the agent can quickly identify vulnerabilities but struggles with exploitation parameters like payload encoding and SQL injection syntax. Network security tasks also yield low performance because the agent disrupts its own network connection by launching intensive scans, highlighting a limit that the agent cannot adapt to the environment it constructed. On real-world tasks, the agent frequently misconfigures Metasploit [metasploit_framework] options and repeatedly attempts exploitation with incorrect settings. This explains the large gap between the success rate of delivery and exploitation.
5.2 Defensive Capability Evaluation
Experiment Setup and Design. In this experiment, we simulate an end-to-end defense process in which the agent is given a software codebase and needs to first identify vulnerabilities and generate corresponding PoCs. It then needs to fix the vulnerabilities and ensure that the generated patches do not affect the codebase’s normal functionality. Since no end-to-end benchmark currently exists, we use two separate benchmarks for this experiment, both derived from real-world codebases in OSS-Fuzz [serebryany2017oss]: CyberGym [wang2025cybergym] for PoC generation and PatchAgent [yu2025patchagent] for patch generation. For PatchAgent [yu2025patchagent], we do not provide agent with functionality tests to better simulate the real-world end-to-end process where only a vulnerable codebase is given. Instead, we supply only the sanitizer report and execution environment, requiring the agent to complete the whole process independently—including root cause analysis, patch generation, and patch validation with its generated test cases. Note that we do not include the agent proposed by PatchAgent in our comparison, as it cannot perform the end-to-end task. We use the same agent scaffold and model as the offensive capability evaluation: OpenHands and Claude Sonnet 4.5 (denoted as OpenHands-Def).
PoC Generation Results. We run a randomly selected subset of 300 cases with OpenHands-Def. It achieves a similar success rate as Claude Sonnet 4.5 with a custom agent scaffold on the whole set reported in its system card, that is, 28.9% [claude45]. By analyzing the agent’s trajectories, we find that the agent exhibits diverse behaviors, including writing scripts to automatically generate and mutate test cases across various input formats, and leveraging external packages to satisfy semantic constraints needed to trigger vulnerabilities. As shown in Figure 5, the agent allocates most of their reasoning steps to browsing and analyzing the codebase to identify potentially vulnerable sections and then determining the specific conditions required to reach and exploit these code paths. This is also one of the most challenging steps for PoC generation; once the analysis is correct, generation becomes straightforward. Interestingly, the agent also produced PoCs that triggered vulnerabilities in supposedly fixed versions, leading to the discovery of new vulnerabilities and revealing incomplete patches in the original fixes. However, this agent lacks program analysis tools for efficient code retrieval and automated security analysis, which are crucial for analyzing large codebases in a more systematic way.
Patch Generation Results. The agent achieves 85% success rate in generating patches that both fix the security vulnerability and maintain original functionality. This strong performance demonstrates the agent’s strong patch generation capabilities after provided with enough information. Notably, over 80% of the ground-truth patches in the benchmark modify functions that appear in the crash stack traces reported by sanitizers, which reduces the difficulty of root cause analysis. In addition, more than 80% of the ground-truth patches affect fewer than 15 lines of code in total (including both additions and deletions), consistent with prior findings [li2017large] that security patches are generally self-contained. As such, the agent typically just directly applies its generated patches without producing any functionality tests. This result indicates that patching security vulnerabilities is relatively straightforward for SOTA agents once the root cause is given. However, the agent still fails in more challenging cases where the root causes are not present in the stack traces (5%), or where the generated patches affect the functionality or have incorrect formats that cause errors (10%). Such cases require deep analysis of the target codebases, which cannot be handled by SOTA agents.
@@ 6 Expert Survey
Expert surveys are a widely used approach for forecasting AI development timelines and risks [grace2024thousands, ai_forecast]. Accordingly, we conduct an expert survey to assess experts’ perceptions of frontier AI’s current impact on cybersecurity and their forecasts of its future impact.
@@ 7.2 Asymmetry-Based Prediction
Asymmetry in cost of failures. A clear asymmetry exists between attackers and defenders in their goals and requirements. Attackers need just one successful exploit, while defenders must guard against all possible threats and vulnerabilities [singer2014cybersecurity]. This fundamental difference severely constrains defenders’ failure tolerance. For AI-driven threat detection systems, both false positives and false negatives carry serious consequences: false positives disrupt normal functionalities, while false negatives leave threats uncovered. Even a single false negative can compromise the entire defense. Similarly, for automatic patching, ensuring accuracy before deploying machine-generated code is essential, raising capability requirements and necessitating additional verification processes that inevitably require human intervention. Compounding this issue, frontier AI’s inherent probabilistic nature introduces errors that further restrict its deployment in critical defense systems.
@@
Different priorities of scalability and reliability. Attackers prioritize scalability, while defenders emphasize reliability. This makes AI a natural choice for attackers, as its primary capability is reducing human involvement and increasing automation. AI lowers barriers for attackers by decreasing requirements for deep expertise, specific target system knowledge, and execution time and effort. Even if an attack fails against one system, it may succeed against another. For example, in large-scale attacks, attackers may simultaneously target thousands or millions of systems (e.g., Mirai Botnet [Mirai]). Conversely, frontier AI models face reliability limitations in robustness, transparency, and generalizability [zhou2024llm, szymanski2025limitations, gu2024survey]. When used for malware detection, users cannot understand why specific software is flagged as malicious or predict when false predictions will occur. This reliability deficit hinders defenders from fully trusting and deploying AI techniques in critical defense systems including decision-making processes. This asymmetry may be magnified in critical physical infrastructure, where attackers exploit these infrastructures and propagate the damage to other systems.
@@ 7.3 Evolutionary Attack-Defense Dynamics
Long-term, however, the evolutionary trajectory may gradually shift toward defenders as several counterfactors emerge. First, defenders could gain priority access to advanced techniques (e.g., OpenAI’s early access for safety researchers [early]). As frontier AI advances, defenders benefit from improved automation, transitioning from manual patch development to AI-generated patches with automated verification. Second, as defenders use AI to systematically discover and patch vulnerabilities, systems become increasingly robust. In particular, by leveraging AI for formal verification and adopting a secure-by-design approach, defenders can build fundamentally more secure systems with provable guarantees. This continuous security improvement creates an environment where finding new vulnerabilities becomes significantly more challenging and costly for attackers. Once defenders achieve stronger system security, attackers must invest substantially more resources to find attack vectors, potentially making many attacks impractical or cost-prohibitive. Lastly, as regulatory frameworks and law enforcement mechanisms mature, attackers can face more severe penalties once caught, substantially raising failure costs and potentially deterring malicious activities.
@@ 8.1 Continuous Marginal Risk Assessment
Build benchmarks and red-teaming approaches for new risks in AI agents. We need to conduct fine-grained risk classification for agents and build specialized benchmarks for each category, given that agents are highly specialized (e.g., web agents, coding agents, and personal assistants). We can also classify risks based on their targets: AI components or symbolic components. Currently, the predominant risk is through indirect prompt injection, attacking AI through symbolic components (tools). However, there are only a few early attacks targeting symbolic components through AI, such as AI-generated exploits against coding agents. This underscores the critical need for thorough risk evaluations of agents. Existing red-teaming for agents only explores limited risks of specific agents (e.g., interpreter abuse of code agents [wan2024cyberseceval], prompt injection of web agents [liao2024eia, evtimov2025wasp], indirect prompt injection [wang2025agentvigil], and taint-style vulnerability [liumake] of various agents). Developing automated red-teaming frameworks for broader agentic risks is necessary, such as assessing multi-agent system risks through game theory or designing new attacks for systems with multi-modal agents.
Continuously ensure quality and realism of benchmarks. If a security benchmark is flawed, agents can “cheat” by completing the task via unintended pathways rather than the intended exploit. Given the complexity of these tasks—often spanning multiple files, services, and hosts—benchmarks must be carefully constructed, validated, and demonstrably solvable. For instance, Transluce noted that several of the Intercode-CTF benchmark tasks are broken or unsolvable [meng2025docent, yang2023language]. Moreover, human expertise remains essential for creating quality environments and data. Research shows that the automated collection of vulnerable code commits without human oversight often yields non-vulnerable or security-unrelated samples [ullah2024llms]. Balancing quality with scale requires combining expert-created seed samples with automated methods [yang2024seccodeplt]. To ensure robust assessment quality, a combination of LLM judges, dynamic evaluation methods, and static program analysis techniques is recommended. Benchmarks should reflect realistic scenarios; real-world testing would be ideal, but raises challenges like verification difficulties and legal/ethical constraints. A practical approach is to replicate production systems in controlled environments with built-in verification mechanisms to assess frontier AI capabilities. Finally, cybersecurity benchmarks must be continuously updated to keep pace with the rapidly evolving threat landscape. Timely updates surface emerging AI-enabled capabilities, allowing practitioners to prioritize mitigations and validate defenses before widespread exploitation.
@@ 8.2 Enhance Cyber Defenses
Enhance attack detection. We first recommend constructing high-quality benchmarks that minimize the critical data and label issues, such as duplicated data, spurious correlations, biased samples, and wrong and missing labels. A recent research [arp2022and] provided a comprehensive summary of these issues. Given the challenges of collecting high-quality real-world attack data, collecting data from controlled environments [beltiukov2023search] and creating synthetic data [wu2023grim] could be efficient ways. Constructing hidden test sets with unseen attacks will help continuously retrain AI-driven defenses; once the currently hidden test sets are used for training, new test sets must be constructed and kept separate from the training data. Such practices have been applied in the natural language field: hiding test sets (e.g., SQuAD [rajpurkar2018know]), updating test sets with recent data (e.g., LiveBench [white2024livebench]), or estimating and resolving training-test overlap [zhang2024language].
@@
Improve the automation of defenses with provable guarantees. We believe frontier AI will significantly advance the scalability and automation of formal verification and call for increased efforts in this direction. First, we can use and fine-tune models to generate more complex proofs and constraints for theorem proving, as well as invariants for complex programs [yang2024formal]. Second, we can use LLMs to generate necessary inputs for formal verification, such as program invariants and function summaries, where enabling verifiable generation with correctness guarantees is crucial. More broadly, secure code generation is a promising direction that can be used to turn unsafe latency code into safe counterparts. Early research has explored strategies like prefix learning [he2023large] and constrained decoding [fu2024constrained]. We suggest focusing more on fundamental improvements in the models or agents, such as RL-based fine-tuning to improve their inherent capability of generating safe code. We can also design customized models and agents to transpile legacy programs into newer and memory-safe languages (e.g., C to Rust [emre2021translating]). Third, AI can also improve constrained solvers. For example, SOTA solvers rely on random/rule-based mechanisms for backtracking and re-selecting verification paths to resolve conflicts. With frontier AI, solvers could learn more efficient strategies for making these critical decisions. Alongside building better AI-enhanced defenses, society must improve overall security posture and actively deploy these new defensive technologies.
@@
Lesson and recommendations based on AIxCC. The AIxCC competition [AIxCC] highlights valuable insights into the current capabilities and limitations of AI-powered cybersecurity systems. The competition required teams to build systems to automatically analyze large codebases and generate PoCs and patches. Most systems in the final still heavily rely on traditional techniques for vulnerability detection and PoC generation. It shows that AI capabilities in these fundamental steps require significant improvement, which aligns with our study. In contrast, patch generation has shown significant evolution with heavy reliance on LLMs, which represent the SOTA in automated code generation. However, these systems still incorporate substantial human knowledge through carefully crafted workflows and prompts. Notably, team Theori [theori_roboduck] adopts a more agentic strategy by using a large number of agents across all steps, enabling them to achieve third place with simpler code structure and reduced human efforts.  These observations align with our recommendations on developing more advanced agents with agent-determined workflows and more program analysis tools, where agent-determined workflows can improve the generalizability and domain-specific tools can improve the agent’s capabilities in security tasks (e.g., static analysis-based context retrieval and fuzzing-based patch validation). In addition, the agentic systems can provide environment, data, and reward signals to continuously improve models’ capabilities in such systems.